OpenStudy (anonymous):

a certain first order reaction is 45% complete in 65s. what are the rate constant and half life for this process? I really want help with learning how to do this type of problem step by step im lost

OpenStudy (aaronq):

So, we'd wanna use the Arrhenius equation: \(\large [A_t]=[A_0]*e^{-kt}\) we'd have to make some assumptions, like that "45% complete" means that 45% of the reactant has been converted into something else. Have you used this equation before?

OpenStudy (anonymous):

so 55% remains therefore ln.55=-kt+ln1 im not sure what to do next alson not sure y the ln 1

OpenStudy (aaronq):

yes, that's right. you also specify the time to find k. Then you use: \(t_{1/2}=\dfrac{ln(2)}{k}\)

OpenStudy (anonymous):

not sure how to read ur notations

OpenStudy (aaronq):

\(ln(0.55)=-k(65~s)+ln(1) \)

OpenStudy (aaronq):

then half-life\(=t_{1/2}=\dfrac{ln(2)}{k}\)

OpenStudy (anonymous):

do u divide -65 to get k

OpenStudy (aaronq):

\(k=-\dfrac{ln(0.55)}{65~s}\)

OpenStudy (aaronq):

if that's what you mean

OpenStudy (anonymous):

im not sure what dfac means

OpenStudy (aaronq):

are you not seeing the \(\LaTeX \)?

OpenStudy (aaronq):

to get k, divide ln(0.55) by -65 sec

OpenStudy (anonymous):

ok cool I got k now explain y I take ln(2)/k

OpenStudy (aaronq):

explain why? it's a graphical result. It just works out because of the exponential model the equation uses.

OpenStudy (anonymous):

well in my notes I have that t1/2=0.693/k

OpenStudy (aaronq):

ln(2)=0.693

OpenStudy (aaronq):

it's the same thing

OpenStudy (anonymous):

ok cool totally didnt know what are the units

OpenStudy (anonymous):

how long does it take for the reaction to go to ?percents

OpenStudy (aaronq):

the units are based on what you found for k, it should be something like "/sec"

OpenStudy (aaronq):

now that you know k, you can find how long it takes for the reaction to hit 50% or whatever

OpenStudy (anonymous):

yes just need a start on the first one is 25%.

OpenStudy (aaronq):

so set it up and i'll check it

OpenStudy (anonymous):

i would im not sure i know how believe me i dont want you to just me the answer but i am lost or unaware of the process

OpenStudy (anonymous):

is it the same way as above

OpenStudy (aaronq):

yeah, it's the same as above except you're solving for t, not k